# Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning

### v1

@@ 4.2 Lower expectile Q-learning with $\lambda$-return
To further improve LEQ for long-horizon tasks, we use $\lambda$-return instead of $1$-step return for Q-learning. $\lambda$-return allows a Q-function and policy to learn from low-bias multi-step returns [28]. Reducing bias in value estimation with $\lambda$-return is especially important on long-horizon tasks where values for nearby states are similar to each other, as illustrated in Figure 2.
We first define $\lambda$-return of a trajectory $\tau$ in timestep $t$, $Q_{t}^{\lambda}(\tau)$, using $N$-step return, $G_{t:t+N}(\tau)$: [^2] Our $\lambda$-return is slightly different from [31, 11] that puts a high weight to the last $N$-step return, $G_{t:H}(\tau)$.
$G_{t:t+N}(\tau)=\sum_{i=0}^{N-1}\gamma^{i}r(\mathbf{s}_{t+i},\mathbf{a}_{t+i})+\gamma^{N}Q_{\phi}(\mathbf{s}_{t+N},\mathbf{a}_{t+N}),$ (8)
$Q_{t}^{\lambda}(\tau)=\frac{1-\lambda}{1-\lambda^{H-t-1}}\sum_{i=1}^{H-t}\lambda^{i-1}G_{t:t+i}(\tau).$ (9)
Then, we can rewrite the Q-learning loss in Equation 4 with $1$-step return to the one with $\lambda$-return:
$\mathcal{L}^{\lambda}_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H-1}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-Q^{\lambda}_{t}(\tau))\right].$ (10)
Figure 2: $\lambda$-return with noisy critic. For long-horizon tasks, value estimates (blue for high values and white for low values) are likely to be flat for the states far from the reward signal shown with the yellow star. Since Q-values are small and similar to each other in this region, Q-values are easily dominated by noise, leading to a wrong learning signal for the policy, as illustrated in the middle row. $\lambda$-return can effectively reduce the critic noise by considering both short-term and long-term value estimates, leading to a better learning signal for the policy.
@@ 4.3 Lower expectile policy learning with $\lambda$-return
For policy optimization, we can use a deterministic policy $\mathbf{a}=\pi_{\theta}(\mathbf{s})$ and update the policy using the deterministic policy gradients similar to DDPG [22]. [^3] LEQ also works with a stochastic policy; but, a deterministic policy is easier to train in an offline setup. Instead of maximizing the immediate Q-value, $Q_{\phi}(\mathbf{s},\mathbf{a})$, we propose to directly maximize the lower expectile of $\lambda$-return, which is a more accurate learning target for a policy, analogous to the conservative critic target in Section 4.2:
$\mathcal{L}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q_{t}^{\lambda}(\tau)\right]\right].$ (11)
However, due to the expectile term in Equation 11, computing the gradient of $\mathcal{L}_{\pi}^{\lambda}(\theta)$ is not trivial. To estimate this gradient, we propose a differentiable surrogate loss, approximating $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q^{\lambda}_{t}(\tau)\right]$ in Equation 11 with $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$:
^$\hat{\mathcal{L}}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\lvert\tau-\mathbbm{1}\left(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)\right)\rvert\cdot Q_{t}^{\lambda}(\tau)\right].$ (12)
Intuitively, this surrogate loss sets a higher weight ($1-\tau$) on a conservative $\lambda$-return estimation (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)$), encouraging a policy to optimize for this conservative $\lambda$-return. On the other hand, an optimistic $\lambda$-return estimation (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})<Q^{\lambda}_{t}(\tau)$) has a less impact to the policy with a smaller weight ($\tau$). Thus, optimizing this surrogate loss leads to the policy maximizing the lower expectile of $\lambda$-return. We provide a proof in Appendix B saying that the proposed surrogate loss is a better approximation of Equation 11 than directly maximizing Q-values, $Q_{\phi}(\mathbf{s},\mathbf{a})$.
@@ 4.4 Expanding dataset with model rollouts
One of the problem of offline RL is that data distribution is limited to the offline dataset $\mathcal{D}_{\text{env}}$. To tackle this problem, we can simulate the current policy inside the model and use the generated trajectories to expand the dataset, similar to prior works [35, 30]. However, the state coverage will be identical when the policy converges, which might lead to catastrophic forgetting.
To prevent this issue, we expand the dataset using the model rollouts from a noisy exploration policy. Specifically, we execute the exploration policy $\pi_{\exp}(\cdot\mid\mathbf{s})$, which simply adds noise $\epsilon\sim N(0,\sigma_{\text{exp}}^{2})$ to the current policy $\pi_{\theta}(\mathbf{s})$, and generate a trajectory of length $R$ ($R=5$ in this paper). We refer this expanded dataset as $\mathcal{D}_{\text{model}}$. Note that we do not use off-policy actions and rewards for critic and policy updates; instead, we generate $H$-step model rollouts starting from $\mathbf{s}\sim\mathcal{D}_{\text{model}}$ and use them for training the policy and Q-function. Thus, we need to store only the states from the rollouts.
@@ 5 Experiments
In this paper, we propose a novel model-based offline RL method with simple and efficient yet accurate conservative value estimation. Through our experiments, we aim to answer the following questions: (1) Can LEQ solve long-horizon tasks? (2) How does LEQ perform in widely used offline RL benchmarks? (3) Which component enables model-based offline RL to learn the AntMaze tasks?
(a) Hopper
(b) Walker2d
(c) HalfCheetah
@@
(d) Umaze
(e) Medium
(f) Large
(g) Ultra
Figure 4: AntMaze tasks.
@@ 5.1 Tasks
To show the strength of LEQ in solving long-horizon tasks, we use the AntMaze tasks, which aims to navigate a $8$-DOF ant robot to the desired goal position, as shown in Figure 4. Specifically, we use umaze, medium, large datasets from D4RL [6], and ultra dataset from Jiang et al. [15]. Moreover, we evaluate our method on MuJoCo locomotion tasks (Figure 4) with dense rewards with D4RL [6] and NeoRL [26] datasets. Please refer to Appendix A for more experimental details.
@@ 5.2 Compared offline RL algorithms
We compare the performance of LEQ with the state-of-the-art offline RL algorithms. Please note that LEQ uses the same hyperparameters across all tasks, except the expectile parameter, $\tau$.
@@
We consider behavioral cloning (BC[25]; TD3+BC [7], which combines BC loss to TD3; CQL [19], which penalizes the actions out of data distribution; and IQL [17], which utilizes expectile regression to estimate the value function. For locomotion tasks, we also compare with EDAC [1], which penalizes the Q-values according to the uncertainty of Q-functions
@@
We consider MOPO [34] and MOBILE [30], which penalize Q-values according to the transition uncertainty and the bellman uncertainty of a world model, respectively; COMBO [35], which combines CQL with MBPO; RAMBO [27], which trains an adversarial world model against the policy; and CBOP [14], which utilizes multi-step returns for critic updates.
5.3 Results on long-horizon AntMaze tasks
As shown in Table 1, LEQ significantly outperforms the prior model-based approaches for all $8$ datasets. LEQ achieves $58.6$ and $60.2$ for antmaze-large-play and antmaze-large-diverse, while the second best method, RAMBO [27], scores only $0.0$ and $2.4$, respectively. We believe these performance gains come from our conservative value estimation, which works more stable than the uncertainty-based penalization of prior works.
Moreover, LEQ even significantly outperforms the model-free approaches in antmaze-umaze, antmaze-large, and antmaze-ultra. Despite its superior performance, LEQ often shows high variance during training, resulting in worse performance on antmaze-medium. Over the course of training, LEQ mostly achieves high success rates, but the evaluation results sometimes drops to $0\%$ as shown in Appendix, Figure 5. We leave the problem of reducing the high variance of our method in certain environments as a future work.
@@ 5.3 Results on long-horizon AntMaze tasks
Model-free Model-based
Dataset BC TD3+BC CQL IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
antmaze-umaze $65.0$ $78.6$ $74.0$ $87.5$ $0.0$ $80.3$ $25.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{94.4}$$\mathbf{\pm 6.3}$
antmaze-umaze-diverse $55.0$ $71.4$ $84.0$ $62.2$ $0.0$ $57.3$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{71.0}$$\mathbf{\pm 12.3}$
antmaze-medium-play $0.0$ $3.0$ $61.2$ $\mathbf{71.2}$ $0.0$ $0.0$ $16.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $58.8$$\pm 33.0$
antmaze-medium-diverse $0.0$ $10.6$ $53.7$ $\mathbf{70.0}$ $0.0$ $0.0$ $23.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $46.2$$\pm 23.2$
antmaze-large-play $0.0$ $0.0$ $15.8$ $39.6$ $0.0$ $0.0$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{58.6}$$\mathbf{\pm 9.1}$
antmaze-large-diverse $0.0$ $0.2$ $14.9$ $47.5$ $0.0$ $0.0$ $2.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{60.2}$$\mathbf{\pm 18.3}$
antmaze-ultra-play -- -- -- $8.3$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{25.8}$$\mathbf{\pm 18.2}$
antmaze-ultra-diverse -- -- -- $15.6$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{55.8}$$\mathbf{\pm 18.3}$
Total w/o antmaze-ultra $120.0$ $163.8$ $303.6$ $354.1$ $0.0$ $137.6$ $67.0$ $0.0$ $0.0$ $\mathbf{388.8}$
Total -- -- -- $378.0$- -- -- $0.0$ $0.0$ $\mathbf{470.4}$
@@
5.4 Results on MuJoCo Gym locomotion tasks
For D4RL MuJoCo Gym tasks in Table 3, LEQ achieves comparable results with the best score of prior works in $6$ out of $12$ tasks. Furthermore, in Table 2, LEQ outperforms most of the prior works in the NeoRL benchmark, especially in the Hopper and Walker2d domains. These results show that LEQ serves as a general offline RL algorithm, not limited to long-horizon tasks.
Similar to antmaze-medium, LEQ also suffers from the high variance problem. During training, LEQ often achieves high performance, but then, suddenly falls back to 00, as shown in Appendix, Figure 5. This is mainly because the learned models sometimes fail to capture failures (e.g. hopper and walker falling off) and predict an optimistic future (e.g. hopper and walker walking forward).
Table 2: NeoRL results. LEQ and IQL results are averaged over $5$ seeds. The results for prior works are reported following Sun et al. [30] and Qin et al. [26]. MOPO is an improved version of MOPO presented in Sun et al. [30]. We highlight the results that are better than $95\%$ of the best score.
@@ 5.4 Results on MuJoCo Gym locomotion tasks
Table 3: D4RL MuJoCo Gym results. Each number is a normalized score averaged over $100$ trials [6]. Our results are averaged over $5$ seeds. The results for prior works are reported following their respective papers. MOPO is an improved version of MOPO, introduced in Sun et al. [30]. We highlight the results that are better than $95\%$ of the best score.

### v2

@@ 4.2 Lower expectile Q-learning with $\lambda$-return
To further improve LEQ, we use $\lambda$-return instead of $1$-step return for Q-learning. $\lambda$-return allows a Q-function and policy to learn from low-bias multi-step returns (Schulman et al., 2016). Using $N$-step returns $G_{t:t+N}(\mathcal{T})=\sum_{i=0}^{N-1}\gamma^{i}r_{t+i}+\gamma^{N}Q_{\phi}(\mathbf{s}_{t+N},\mathbf{a}_{t+N})$, we define $\lambda$-return of an $H$-step trajectory $\mathcal{T}$ in timestep $t$, $Q_{t}^{\lambda}(\mathcal{T})$ as: [^2] Our $\lambda$-return slightly differs from (Sutton, 1988) that puts a high weight to the last $N$-step return, $G_{t:H}(\mathcal{T})$.
$Q_{t}^{\lambda}(\mathcal{T})=\frac{1-\lambda}{1-\lambda^{H-t-1}}\sum_{i=1}^{H-t}\lambda^{i-1}G_{t:t+i}(\mathcal{T}).$ (8)
Then, we can rewrite the Q-learning loss in Equation 4 with $\lambda$-return targets:
$\mathcal{L}^{\lambda}_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H-1}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-Q^{\lambda}_{t}(\mathcal{T}))\right].$ (9)
@@ 4.3 Lower expectile policy learning with $\lambda$-return
For policy optimization, we use a deterministic policy $\mathbf{a}=\pi_{\theta}(\mathbf{s})$ and update the policy using the deterministic policy gradients similar to DDPG (Lillicrap et al., 2016). [^3] LEQ also works with a stochastic policy; but, a deterministic policy is sufficient for our experiments. To provide more accurate learning targets for the policy, instead of maximizing the immediate Q-value, $Q_{\phi}(\mathbf{s},\mathbf{a})$, we maximize the lower expectile of $\lambda$-return, analogous to our conservative critic learning in Section 4.2:
$\mathcal{L}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}}}\left[\sum_{t=0}^{H}\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[Q_{t}^{\lambda}(\mathcal{T})\right]\right].$ (10)
However, because of the expectile term, $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]$, we cannot directly compute the gradient of $\mathcal{L}_{\pi}^{\lambda}(\theta)$. To change the expectile to expectation, we use the relationship $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]=\frac{\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\cdot Q_{t}^{\lambda}\right]}{\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\right]}$, and optimize the unnormalized version, $\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\cdot Q_{t}^{\lambda}\right]$. By approximating $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]$ with the learned Q-estimator $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$, we derive a differentiable surrogate loss of Equation 10:
^$\hat{\mathcal{L}}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\lvert\tau-\mathbbm{1}\left(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T})\right)\rvert\cdot Q_{t}^{\lambda}(\mathcal{T})\right].$ (11)
Intuitively, this surrogate loss sets a higher weight ($1-\tau$) on a conservative $\lambda$-return estimates (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T})$), encouraging a policy to optimize for this conservative $\lambda$-return. On the other hand, an optimistic $\lambda$-return estimates (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})<Q^{\lambda}_{t}(\mathcal{T})$) has a less impact to the policy with a smaller weight ($\tau$). We provide a proof in Appendix B showing that the proposed surrogate loss provides a better approximation of Equation 10 than the immediate Q-value, $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
@@ 4.4 Expanding dataset with model rollouts
One of the problems of offline RL is that data distribution is limited to the offline dataset $\mathcal{D}_{\text{env}}$. To tackle this problem, we expand the dataset using simulated trajectories, which we refer to as $\mathcal{D}_{\text{model}}$ (Yu et al., 2021). To diversify the simulated trajectories in $\mathcal{D}_{\text{model}}$, we use a noisy exploration policy, which adds a noise $\epsilon\sim N(0,\sigma_{\text{exp}}^{2})$, to the current policy and generate a trajectory of length $R$.
@@ 5 Experiments
In this paper, we propose a novel model-based offline RL method with simple, efficient, yet accurate conservative value estimation. Through our experiments, we aim to answer the following questions: (1) Can LEQ solve diverse domains of problems, including both dense-reward tasks and long-horizon sparse-reward tasks? (2) Can LEQ be applied to pixel-based environments? (3) How individual components of LEQ affect the performance?
(a) Umaze
(b) Medium
(c) Large
(d) Ultra
Figure 3: AntMaze tasks.
(e) Hopper
(f) Walker2d
(g) HalfCheetah
@@ 5.1 Tasks
To verify the strength of our low-bias model-based conservative value estimation in diverse domains, we test LEQ on four benchmarks: D4RL AntMaze, D4RL MuJoCo Gym (Fu et al., 2020), NeoRL (Qin et al., 2022), and V-D4RL (Lu et al., 2023). We first test on long-horizon AntMaze tasks: umaze, medium, large from D4RL, and ultra from Jiang et al. (2023), as shown in Figure 4. We also evaluate LEQ on locomotion tasks (Figure 4): state-based tasks from D4RL, NeoRL and pixel-based tasks from V-D4RL. Please refer to Appendix A for more experimental details.
@@ 5.2 Compared offline RL algorithms
We consider behavioral cloning (BC(Pomerleau, 1989); TD3+BC (Fujimoto and Gu, 2021), which combines BC loss to TD3; CQL (Kumar et al., 2020), which penalizes out-of-distribution actions; and IQL (Kostrikov et al., 2022), which uses upper-expectile regression to estimate the value function. For locomotion tasks, we also compare with EDAC (An et al., 2021), which penalizes Q-values based on its uncertainty. For pixel-based tasks, we compare with DrQ+BC (Lu et al., 2023), which combines BC loss to DrQ-v2 (Yarats et al., 2022); ACRO, which learns representations with a multi-step inverse dynamics model.
@@
We consider MOPO (Yu et al., 2020) and MOBILE (Sun et al., 2023), which penalize Q-values according to the transition uncertainty and the Bellman uncertainty of a world model, respectively; COMBO (Yu et al., 2021), which combines CQL with MBPO; RAMBO (Rigter et al., 2022), which trains an adversarial world model against the policy; and CBOP (Jeong et al., 2023), which utilizes multi-step returns for critic updates; IQL-TD-MPC (Chitnis et al., 2024), which extends TD-MPC (Hansen et al., 2022) to offline setting with IQL. For pixel-based environments, we consider OfflineDV2 (Lu et al., 2023), which penalizes Q-values according to the dynamics errors, and ROSMO (Liu et al., 2023), which uses one-step model rollouts for policy improvement. LEQ follows MOBILE for most implementation details but implemented in JAX (Bradbury et al., 2018), which makes it $6$ times faster than the PyTorch versions of MOBILE and CBOP. Please refer to Table 8 for detailed comparison.
Sequence modeling for offline RL.
We consider TT (Janner et al., 2021), which trains a Transformer model (Vaswani, 2017) to predict offline trajectories and applies beam search to find the best trajectory; and TAP (Jiang et al., 2023), which improves TT by quantizing the action space with VQ-VAE (Van Den Oord et al., 2017).
@@
Model-free Seq. modeling Model-based
Dataset CQL IQL TT TAP COMBO RAMBO MOBILE CBOP IQL-TD-MPC LEQ (ours)
umaze $74.0$ $87.5$ $\mathbf{100.0}$ $81.5$ $80.3$ $25.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $52.0$ $94.4$$\pm 6.3$
umaze-diverse $\mathbf{84.0}$ $62.2$ $21.5$ $68.5$ $57.3$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $72.6$ $71.0$$\pm 12.3$
medium-play $61.2$ $71.2$ $\mathbf{93.3}$ $78.0$ $0.0$ $16.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{88.8}$ $58.8$$\pm 33.0$
medium-diverse $53.7$ $\mathbf{70.0}$ $\mathbf{100.0}$ $85.0$ $0.0$ $23.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $40.3$ $46.2$$\pm 23.2$
large-play $15.8$ $39.6$ $66.7$ $\mathbf{74.0}$ $0.0$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $66.6$ $58.6$$\pm 9.1$
large-diverse $14.9$ $47.5$ $60.0$ $\mathbf{82.0}$ $0.0$ $2.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $4.0$ $60.2$$\pm 18.3$
ultra-play -- $8.3$ $20.0$ $22.0$- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $20.6$ $\mathbf{25.8}$$\mathbf{\pm 18.2}$
ultra-diverse -- $15.6$ $33.3$ $26.0$- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $3.6$ $\mathbf{55.8}$$\mathbf{\pm 18.3}$
Total w/o ultra $303.6$ $354.1$ $441.5$ $\mathbf{469.0}$ $137.6$ $67.0$ $0.0$ $0.0$ $324.3$ $388.8$
Total -- $378.0$ $\mathbf{494.8}$ $\mathbf{517.0}$- -- $0.0$ $0.0$ $348.5$ $470.4$
@@
Table 2: NeoRL results. LEQ and IQL results are averaged over $5$ seeds. The results for prior works are reported following Sun et al. (2023) and Qin et al. (2022). MOPO is an improved version of MOPO presented in Sun et al. (2023). We highlight the results better than $95\%$ of the best score.
@@
5.3 Results on long-horizon AntMaze tasks
As shown in Table 1, LEQ significantly outperforms the prior model-based approaches. For example, LEQ achieves $60.2$ and $55.8$ for large-diverse and ultra-diverse, while the second best method, IQL-TD-MPC (Chitnis et al., 2024), scores only $4.0$ and $3.6$, respectively. We believe these performance gains come from our conservative value estimation, which works more stable than the uncertainty-based penalization of prior works. Moreover, LEQ even significantly outperforms the model-free approaches in umaze, large, and ultra mazes, and outperforms sequence modeling methods, TT and TAP, which serve as strong baselines for AntMaze tasks, in the most challenging ultra mazes, showing the advantage of utilizing low-bias multi-step return on long-horizon tasks.
Figure 5: Failure in medium mazes. The agent plans impossible trajectories on certain states (red circles).
Despite its superior performance, LEQ shows high variance in the performance on antmaze-medium. We found that the medium mazes include states separated by walls that are very close to each other (denoted as red circles in Figure 5), such that all of the learned world models falsely believe the agent can pass through the walls. This incorrect prediction makes the agent to plan faster, but impossible trajectories as shown in Figure 5. We believe that this could be addressed by employing improved world models or increasing the number of ensembles for the world models.
Table 3: D4RL MuJoCo Gym results. Each number is a normalized score averaged over $100$ trials (Fu et al., 2020). Our results are averaged over $5$ seeds. The results for prior works are reported following their respective papers. MOPO is an improved version of MOPO, introduced in Sun et al. (2023). We highlight the results that are better than $95\%$ of the best score.
Model-free Seq. modeling Model-based
Dataset CQL EDAC IQL TT TAP MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
hopper-r $5.3$ $25.3$ $7.6$ $6.9$ $31.7$ $17.9$ $25.4$ $\mathbf{31.9}$ $\mathbf{32.8}$ $\mathbf{32.4}$$\pm 0.3$
hopper-m $61.9$ $\mathbf{101.6}$ $66.3$ $67.4$ $63.4$ $62.8$ $97.2$ $87.0$ $\mathbf{106.6}$ $\mathbf{102.6}$ $\mathbf{103.4}$$\pm 0.3$
hopper-mr $86.3$ $\mathbf{101.0}$ $94.7$ $99.4$ $87.3$ $\mathbf{99.4}$ $\mathbf{103.5}$ $89.5$ $99.5$ $\mathbf{104.3}$ $\mathbf{103.9}$$\pm 1.3$
hopper-me $96.9$ $\mathbf{110.7}$ $91.5$ $\mathbf{110.0}$ $105.5$ $81.6$ $\mathbf{111.1}$ $88.2$ $\mathbf{112.6}$ $\mathbf{111.6}$ $\mathbf{109.4}$$\pm 1.8$
walker2d-r $5.4$ $16.6$ $5.2$ $5.9$ $7.4$ $7.0$ $0.0$ $17.9$ $17.8$ $\mathbf{21.5}$$\pm 0.1$
walker2d-m $79.5$ $\mathbf{92.5}$ $78.3$ $84.9$ $64.9$ $81.3$ $84.1$ $81.9$ $84.9$ $\mathbf{87.7}$ $74.9$$\pm 26.9$
walker2d-mr $76.8$ $87.1$ $73.9$ $89.2$ $66.8$ $85.6$ $56.0$ $89.2$ $89.9$ $92.7$ $\mathbf{98.7}$$\pm 6.0$
walker2d-me $109.1$ $\mathbf{114.7}$ $109.6$ $101.9$ $107.4$ $\mathbf{112.9}$ $103.3$ $56.7$ $\mathbf{115.2}$ $\mathbf{117.2}$ $108.2$$\pm 1.3$
halfcheetah-r $31.3$ $28.4$ $11.8$ $6.1$ $\mathbf{38.5}$ $\mathbf{38.8}$ $\mathbf{39.5}$ $\mathbf{39.3}$ $32.8$ $30.8$$\pm 3.3$
halfcheetah-m $46.9$ $65.9$ $47.4$ $46.9$ $45.0$ $73.0$ $54.2$ $\mathbf{77.9}$ $\mathbf{74.6}$ $\mathbf{74.3}$ $71.7$$\pm 4.4$
halfcheetah-mr $45.3$ $61.3$ $44.2$ $44.1$ $40.8$ $\mathbf{72.1}$ $55.1$ $\mathbf{68.7}$ $\mathbf{71.7}$ $66.4$ $65.5$$\pm 1.1$
halfcheetah-me $95.0$ $\mathbf{106.3}$ $86.7$ $95.0$ $91.8$ $90.8$ $90.0$ $95.4$ $\mathbf{108.2}$ $\mathbf{105.4}$ $\mathbf{102.8}$$\pm 0.4$
Total $739.7$ $911.4$ $717.2$ $747.5$ $844.0$ $802.0$ $812.4$ $\mathbf{959.5}$ $\mathbf{953.4}$ $\mathbf{923.2}$
5.4 Results on MuJoCo Gym locomotion tasks
For the NeoRL benchmark in Table 2, LEQ outperforms most of the prior works, especially in the Hopper and Walker2d domains. Furthermore, for D4RL MuJoCo Gym tasks in Table 3, LEQ achieves comparable results with the best score of prior works in $7$ out of $12$ tasks, These results show that LEQ serves as a general offline RL algorithm, widely applicable to various domains.
Figure 6: High variance during training. Our algorithm experiences oscillation due to optimistic imaginations near the initial states.
Table 4: V-D4RL results. We report the mean and standard deviation of returns over $3$ seeds. For ROSMO, we replace the categorical distribution in their official code with Gaussian distribution to support continuous action spaces. For MOPO, we implement MOPO on top of DreamerV3. The results for other prior works are reported following Lu et al. (2023); Islam et al. (2023).